If carbon-carbon Bond length in diamond is 1.5 angstorm then what will be the edge length of the cubic closed packed unit cell of diamond having alternate tetrahedral voids occupied?
1
Expert's answer
2019-05-27T06:09:48-0400

Given:

"r_{c-c} = 1.5 { \ \ Angstrom }"

Solution:

CCP cell:



From triangle ABC:



"AB= r_{c-c}""AC= \sqrt{\frac{(AB)^2}{2} }"

Edge length is:


"L = 2 * AC = 2 * \sqrt{\frac{(AC)^2}{2}} = \sqrt{2(AC)^2} = \sqrt{2r_{c-c}^2}"

Calculating:


"L =\sqrt{2r_{c-c}^2} = \sqrt{2*1.5^2}\approx 2.12 \ Angstrom"

Answer: 2.12 Angstrom

References:

  • https://hemantmore.org/chemistry-1/c11201018/1899/